OpenStudy (simplymarie_x):

Give an example of an expression with a rational exponent and explain how to convert it into radical form.

OpenStudy (solomonzelman):

The general rule is: \(\tiny \\[0.5em]\) \(\color{black}{\large \displaystyle x^{ \color{red}{ a}/\color{blue}{b}}=\sqrt[\color{blue}{b}]{x^{\color{red}{ a}}} }\)

OpenStudy (solomonzelman):

For example, \(\tiny \\[0.8em]\) \(\color{black}{\large \displaystyle x^{4/7}=\sqrt[7]{\,x^4\,} }\)

OpenStudy (solomonzelman):

Or, another example, \(\color{black}{\large \displaystyle x^{1/2}=\sqrt[2]{x^1}=\sqrt{x}}\) ------------------------------ Remark: \(\color{black}{ \sqrt[2]{f(x)}}\) is the same thing as \(\color{black}{ \sqrt{f(x)}}\), and \(\color{black}{ \displaystyle x^1=x}\), this is how I wrote my 3rd (final) step. ------------------------------

OpenStudy (solomonzelman):

It doesn't have to be \(x\), it can be a number as well. For example, \(\tiny \\[0.8em]\) \(\color{black}{\large \displaystyle 7^{2/3}=\sqrt[3]{\,7^2\,} }\).

OpenStudy (solomonzelman):

I was wondering what they mean in politics by "Radical Muslims" ? Is it \(\color{black}{\large \displaystyle \sqrt{\normalsize\rm Muslims}}\) ?

OpenStudy (solomonzelman):

Just kidding ... Anyway, do you understand the rule? Do you have any questions?